In each of the following functions determine it's





A.domain





B.range





C.x-intercept





D.zeroes





E.vertical asymptote





1y=log5(x-1)+3





2.y=log(3x-2)

1
Expert's answer
2021-11-12T08:10:48-0500

It is Assumed the logarithm given are to base 10.


"y=log5(x-1)+3"


it's Domain


For the log to be positive "x>1"

"(1,\infin)"


Range


All real numbers

"(-\infin,\infin)"


x-intercept

log "5(x-1)+3=0"

"log5(x-1)=-3"

"5x-5=0.001"

"x=1.0002"


"(1.0002,0)"


Zeros


At "y=0"

"x=1.0002"



Vertical Asymptote

For "5(x-1)>0"

"x>1"

Asymptote is "x=1"


"y=log(3x-2)"


Domain

For log to be positive "3x-2>0\implies\>x>\frac{2}{3}"

"(\frac{2}{3},\infin)"


Range


All real values

"(-\infin,\infin)"


x-intercept

log "(3x-2)=0"

"3x-2=1"

"x=1"

"(1,0)"


Zeros

For "y=0,\quad x=1"


Vertical Asymptote

"3x-2>0"

Vertical Asymptote is "x=\frac{2}{3}"